OpenStudy (anonymous):

Let T:Rn->Rn be an invertible linear transformation, and let S and U be functions from Rn into Rn such that S(T(x))=x and U(T(x))=x for all x in Rn. Show that U(v)=S(v) for all v in Rn. This will show that T has a unique inverse.

OpenStudy (anonymous):

We are going to do a proof by contradiction. Assume that \[S\ne U.\]Then there must exist at least one vector v in Rn such that:\[S(v)\ne U(v).\] However, since T is an invertible linear transformation, there must exist a v' in Rn such that:\[T(v')=v.\]Therefore:\[S(v)=S(T(v'))=v'=U(T(v'))=U(v)\]which is a contradiction to S(v) not being equal to U(v). Thus my initial statement was false, and U=S.

OpenStudy (anonymous):

for all x in Rn S(T(x))=U(T(x)) then SoT = UoT........let H be an inverse of T ( ToH = Id ) (right inverse ) Then SoToH = UoToH ====> So Id = Uo Id====> S=U